A grocery clerk sets up a display of 12-pack cartons of cola. There are 15 cartons at the base of the triangle and one at the top. How many cartons of cola are needed for the complete display?
A.180
B.30
C.120
D.15

Answer by stanbon(75887)

A grocery clerk sets up a display of 12-pack cartons of cola. There are 15 cartons at the base of the triangle and one at the top. How many cartons of cola are needed for the complete display?
1+2+3+...+13+14+15
= (15/2)(1+15) = 15*8 = 120
=================================
cheers,
Stan H.
====================
A.180
B.30
C.120
D.15